# Multimodal Fact Checking with Unified Visual, Textual, and Contextual Representations

### v1

@@ 2 Related Work
Recent research in automated fact-checking has highlighted the growing importance of incorporating multiple modalities, to tackle the diverse and evolving forms of misinformation [22]. Early works, such as FEVER [23] and CLEF2018 [24] are primarily focused on verifying textual claims, laying foundational methods for claim verification based solely on textual evidence. However, later studies found that misinformation exploits images, videos, and audio alongside text to build convincing narratives [25, 26]. These studies have revealed the limitations of purely text-based fact-checking methods and sparked a shift toward multimodal fact-checking. To tackle the limitation of text-based methods, systems were designed to jointly process and reason over diverse types of content. For example, several studies, such as [16] and [27, 28], have explored architectures that integrated textual and visual features. These models employ mechanisms like attention or contrastive learning to enhance detection accuracy. Recent work by [29] shows that a probing classifier combining separate text and image embeddings can outperform intrinsic VLM features on datasets like Factify 2. In addition to these developments, comprehensive surveys, such as [10], offers detailed overview regarding the emerging field of multimodal fact-checking. They highlighted both the technical challenges and the promising research directions ahead. Key challenges include aligning information across modalities, managing incomplete or noisy evidence, and ensuring scalability for practical deployment.
Despite significant progress, effectively integrating multimodal information remains an open research problem. This challenge continues to motivate the development of new architectures and learning methods. Robust fact verification in multimodal contexts still requires innovative solutions.
@@ 4 Methodology
In this section, we have reported our proposed framework, schematically depicted in Figure 4. Our framework has four components, i.e. (i) text module, (ii) image module, (iii) fusion module, and (iv) classification module. Each component of our framework is described as follows,
@@ 5 Experiments:
In this section, we have presented the experiments to evaluate the effectiveness of our framework. We have assessed the impact of different learning strategies, model architectures, and training choices, with a focus on how fusion methods and element-wise operations enhance multimodal relationships.
@@ 5.1 Experimental setup:
All experiments were conducted on a single NVIDIA A100 80GB GPU using PyTorch and Hugging Face Transformers. We evaluated model performance using the weighted F1 score, which balances class-wise F1 scores based on support (i.e., true instance count per class). This metric provides a fairer summary under label imbalance compared to unweighted F1. Formally, if $F_{i}$ is the F1 score and $n_{i}$ is the support for class $i$, then:
@@ 5.3 Experimental Variants:
Without Contrastive Head: In this version, the contrastive projection head and the associated loss are omitted. The model relies solely on cross-entropy loss applied to the fused multimodal representation for classification.
The only difference is the inclusion of contrastive supervision. Results are presented in Tables 4, 5, with discussion in Section 6.
@@ 6 Results and Discussion:
We evaluated our proposed framework on the Factify 2 dataset using multiple combinations of language and vision backbones. Performance is measured using weighted F1 scores across five fine-grained veracity labels. We have reported both overall and class-wise results to assess the model’s capabilities. As shown in Tables 4 and 5, all model variants significantly outperform the baseline. Additionally, models integrated with contrastive head consistently outperform their non-contrastive counterparts. This trend holds across the visual encoders, confirming the robustness of our architecture. Notably, the marginal gains are observed in “Insufficient” and “Refute” categories that require resolving minor differences between text and images. These results show that our fusion method and contrastive learning help the model to better connect text and images.
@@ 7 Error analysis:
Qualitative Analysis: To better understand model behavior, we examined the mismatches between our approach and the baseline. As shown in Figure 5, the baseline overlooked the visual modality entirely and based its decision on text alone, predicting Insufficient Text. In contrast, our model incorporated the OCR-detected credit “Helen Sloan/HBO” from the image, identifying it as a licensed promotional photograph a detail suggesting the image does not contribute new factual content. By recognizing the lack of substantive support in both text and image, our model rightly predicted Insufficient Multimodal. This highlights how image provenance, even in OCR form, enhances factual reasoning. Additional examples are discussed below:
@@
We have presented detailed examples mentioned by the original ID of the samples as per the dataset, comparing predictions from our approach against the baseline. These examples illustrate how OCR-based information like photographer credits, agency marks, or image overlays can either resolve or confuse multimodal evidence verification.
Figure 5: Example of qualitative analysis, Sample from the dataset.
@@
  • ID 4681: The baseline model, using only textual signals, interpreted the phrase “Made in India vaccines” as concrete evidence supporting the claim thus predicting Support_Text. However, the OCR-extracted string “COLO STORAE” (interpreted as “cold storage”) hinted at logistical or infrastructure gaps in the vaccine delivery, casting doubt on the sufficiency of the claim’s evidence. Our model’s multimodal fusion recognized this ambiguity and opted for Insufficient_Text, aligning with the ground truth. This example shows how even noisy OCR text can surface hidden context missed by text-only models.
  • ID 6968: The baseline inferred support based on name and context matches in the text (e.g., “Governor Jagdeep Dhankar”), resulting in a Support_Text prediction. However, the image’s OCR output “ANI” indicated it was a generic press photo from a news agency, lacking evidentiary value. Our model correctly interpreted this, combining the weak textual alignment with the non-informative image tag to determine that neither modality offered enough proof thus choosing Insufficient_Multimodal. This highlights the model’s ability to ignore superficial visual cues.
@@
  • ID 7171: Here, our model was overly cautious. The OCR tag “ANI” led it to interpret the image as a generic stock photo, causing it to discount visual evidence. It then judged the claim as lacking visual proof and labeled it Insufficient_Multimodal. However, the textual portion “will lay foundation stone” is a direct and verifiable event announcement, and the image (even if generic) serves as a credible contextual anchor. The baseline, focusing on the assertive language in the text, correctly predicted Support_Multimodal. This example shows that OCR can occasionally mislead, especially when image content is generic but still contextually supportive.
@@ 8 Conclusion:
This paper introduces a unified framework for fine-grained multimodal fact-checking that jointly reasons over textual and visual evidence. Our architecture integrates structured representations from pre-trained language and vision models using a relational fusion module. It further employs a contrastive learning objective to enhance cross-modal alignment. This design allows the model to better capture fine agreements and contradictions between claims and evidence across modalities. As our approach outperforms the baseline, across multiple configurations. It achieves particularly notable gains in complex classes such as "Insufficient" and "Refute", where multimodal reasoning is critical. Statistical tests and qualitative analyses confirm that the improvements are consistent and meaningful. These gains are systematic, rather than incidental. Overall, our results emphasize the importance of explicit cross-modal alignment and representation learning in advancing automated fact verification.
@@ 9 Limitation:
  • Dependence on OCR quality: The model incorporates OCR-extracted text from images, which can vary widely in accuracy and relevance. In cases of noisy or misleading OCR outputs, the model may misclassify due to spurious visual-textual alignment.
  • No evidence retrieval component: Our framework assumes that relevant evidence both textual and visual is already provided. It does not includes any retrieval pipeline to source additional or more reliable evidences from external knowledge bases or web sources.
  • Limited visual understanding: Although image features are included via pre-trained encoders, the model lacks deeper visual reasoning capabilities such as object detection, scene understanding, or temporal cues that could improve evidence grounding.
  • Restricted modalities: The current system handles only text and image modalities. It does not address audio, video, or temporal multimodal misinformation, which are common in modern social media content.

### v2

@@ 2 Related Work
Recent research in automated fact-checking has highlighted the growing importance of incorporating multiple modalities, to tackle the diverse and evolving forms of misinformation [22]. Early works, such as FEVER [23] and CLEF2018 [24] are primarily focused on verifying textual claims, laying foundational methods for claim verification based solely on textual evidence. However, later studies found that misinformation exploits images, videos, and audio alongside text to build convincing narratives [25, 26]. These studies have revealed the limitations of purely text-based fact-checking methods and sparked a shift toward multimodal fact-checking. To tackle the limitation of text-based methods, systems were designed to jointly process and reason over diverse types of content. For example, several studies, such as [16] and [27, 28], have explored architectures that integrated textual and visual features. These models employ mechanisms like attention or contrastive learning to enhance detection accuracy. Recent work by [29] shows that a probing classifier combining separate text and image embeddings can outperform intrinsic VLM features on datasets like Factify 2. In addition to these developments, comprehensive surveys, such as [10], offers detailed overview regarding the emerging field of multimodal fact-checking. They highlighted both the technical challenges and the promising research directions ahead. Key challenges include aligning information across modalities, managing incomplete or noisy evidence, and ensuring scalability for practical deployment. Despite significant progress, effectively integrating multimodal information remains an open research problem. This challenge continues to motivate the development of new architectures and learning methods. Robust fact verification in multimodal contexts still requires innovative solutions.
@@ 4 Methodology
In this section, we have reported our proposed framework, schematically depicted in Figure 4. Our framework has four components, i.e. (i) text module, (ii) image module, (iii) fusion module, and (iv) classification module. Each component of our framework is described as follows:
@@ 5 Experiments:
In this section, we have presented the experiments to evaluate the effectiveness of our framework. We have assessed how learning strategies, architectures, and fusion methods with element-wise operations affect multimodal performance.
@@ 5.1 Experimental setup:
All experiments were conducted on a single NVIDIA A100 80GB GPU using PyTorch and Hugging Face Transformers. We have evaluated model performance using the weighted F1 score, which balances class-wise F1 scores based on support (i.e., true instance count per class). This metric provides a fairer summary under label imbalance compared to unweighted F1. Formally, if $F_{i}$ is the F1 score and $n_{i}$ is the support for class $i$, then:
@@ 5.3 Experimental Variants:
Without Contrastive Head: In this version, the contrastive projection head and the associated loss are omitted. The model relies solely on cross-entropy loss applied to the fused multimodal representation for classification. Results are presented in Tables 4, 5, with discussion in Section 6.
@@ 6 Results and Discussion:
We have evaluated our proposed framework on Factify 2 using various language and vision encoder backbones. Performance is measured via weighted F1 across five veracity labels. As shown in Tables 4 and 5, all variants significantly outperforms the baseline, models integrated with contrastive-head consistently outperform their non-contrastive counterparts. Specifically gains are most notable in “Insufficient” and “Refute” cases. This highlights the effectiveness of our fusion and contrastive learning objective in aligning text and image evidence.
@@ 7 Error analysis:
Qualitative Analysis: To better understand model behavior, we analyzed the mismatches between our approach and the baseline. As shown in Figure 6,the baseline relied only on text, predicting Insufficient Text. However, Our model used the OCR credit “Helen Sloan/HBO” to recognize the image as a licensed promotional photograph, detail suggesting the image does not contribute new factual content, and correctly predicting Insufficient Multimodal. By recognizing the lack of substantive support in both text and image, our model rightly predicted Insufficient Multimodal. This highlights the role of OCR-extracted image sources in improving factual reasoning. Additional examples are discussed below:
@@
Here, we have presented the dataset-referenced examples comparing our model with the baseline, showing how OCR cues such as credits, agency marks, or overlays can clarify or complicate the multimodal verification.
@@
  • ID 4681: The baseline, relying only on text, misclassified “Made in India vaccines” as Support_Text. In contrast, our model’s multimodal fusion leveraged the OCR cue “COLO STORAE” (cold storage) to detect ambiguity and correctly predicted Insufficient_Text. This illustrates how even noisy OCR can reveal context missed by text-only models.
  • ID 6968: The baseline predicted Support_Text based on name and context matches in text (e.g., “Governor Jagdeep Dhankar”). In contrast, our model used the OCR cue “ANI” to identify the image as a generic press photo and, combining it with weak textual alignment, correctly predicted Insufficient_Multimodal. This demonstrates its ability to dismiss superficial visual signals.
@@
  • ID 7171: In this case, our model was overly cautious. The OCR tag “ANI” led it to dismiss the image as generic, resulting in an Insufficient_Multimodal prediction. However, the textual portion “will lay foundation stone” is a direct and verifiable event announcement, and the image provided valid context. The baseline correctly predicted Support_Multimodal, highlighting that OCR cues can sometimes mislead when visual content, though generic, remains contextually relevant.
    (a) Roberta+ResNet50 (with contrastive head)
    (b) Roberta+ResNet50 (with contrastive head)
    (c) Roberta+ResNet50 (without contrastive head)
    (d) Roberta+ResNet50 (without contrastive head)
    (e) Roberta + ViT (with contrastive head)
    (f) Roberta + ViT (with contrastive head)
    (g) Roberta + ViT (without contrastive head)
    (h) Roberta + ViT (without contrastive head)
    (i) DeBERTa + ViT (with contrastive head)
    (j) DeBERTa + ViT (with contrastive head)
    (k) DeBERTa + ViT (without contrastive head)
    (l) DeBERTa + ViT (without contrastive head)
    (m) DeBERTa + ResNet50 (with contrastive head)
    (n) DeBERTa + ResNet50 (with contrastive head)
    (o) DeBERTa + ResNet50 (without contrastive head)
    (p) DeBERTa + ResNet50 (without contrastive head)
    (q) SBERT + ResNet50 (with contrastive head)
    (r) SBERT + ResNet50 (with contrastive head)
    (s) SBERT + ResNet50 (without contrastive head)
    (t) SBERT + ResNet50 (without contrastive head)
    Figure 5: Comparison of discordant-pair barplots (first column per model) and McNemar 2×2 contingency heatmaps (second column per model), both with and without the contrastive head. Models (top to bottom): Roberta + ResNet50, Roberta + ViT, DeBERTa + ViT, DeBERTa + ResNet50, SBERT + ResNet50.
@@
Figure 6: Example of qualitative analysis, Sample from the dataset.
@@ 8 Conclusion:
This paper presents a unified framework for fine-grained multimodal fact-checking that jointly reasons over text and images. It integrates structured representations from pre-trained encoders through a relational fusion module and enhances cross-modal alignment via contrastive learning. This enables the model to capture subtle agreements and contradictions between claims and evidence. Our approach consistently outperforms the baseline, with significant gains in complex classes like "Insufficient" and "Refute". Statistical and qualitative analyses confirms that these improvements are consistent and non-random. Overall, these results highlights the value of explicit cross-modal alignment and representation learning for reliable fact verification.
@@ 9 Limitation:
  • Dependence on OCR quality: The model relies on OCR-extracted text, which can vary in quality. Noisy or misleading OCR may cause incorrect predictions due to false visual-textual alignment.
  • No evidence retrieval component: Our framework assumes relevant text and image evidence is already provided. It does not include a retrieval mechanism to source evidence from external knowledge bases or the web.
  • Limited visual understanding: Although pre-trained vision encoders are used, the model lacks deeper visual reasoning, such as object detection, scene understanding, or temporal interpretation.
  • Restricted modalities: The system currently supports only text and image. It does not handle audio, video, or temporal misinformation, which are increasingly common in real-world settings.